find the distance b/w the points A(a,b) and B(b,a) if     a-b = 4 

Grade:10

1 Answers

Abu Azbar
6 years ago
A(a,b) and B(b,a)AB = √{a-b}^2+{b-a}^2= √(4)^2+(-4)^2=√(16+16)=√(32)Therefore the distance between points A and B is √32